OpenStudy (anonymous):

solve for x ln(8x-1)=2

OpenStudy (anonymous):

ln(8x -1) = 2 8x -1 = e^2 8x = e^2 +1 Divide by 8 x = 1/8(e^2 + 1)

OpenStudy (anonymous):

log(8 x-1) = 2 Cancel logarithms by taking exp of both sides: 8 x-1 = e^2 Add 1 to both sides: 8 x = 1+e^2 Divide both sides by 8: x = 1/8 (1+e^2) Now test that this solution is appropriate by substitution into the original equation: Check the solution x = 1/8 (1+e^2): log(8 x-1) => log(-1+8/8 (1+e^2)) = 2 So the solution is correct. Thus, the solution is: x = 1/8 (1+e^2)